OpenStudy (anonymous):

Amy sold pens and pencils for a fundraiser at school. Pencils sold for $1.50, and pens sold for $2. She sold a total of 35 items and made $61.50. How many pens did Amy sell?

OpenStudy (anonymous):

p + i =35 1.5 p + 2i = 61.50 1.5 (35-i) +2i =61.50 52.5 - 1.5i + 2i =61.50 .5i = 9.00 i = 18 So 18 pens where sold.